OpenStudy (anonymous):

2x + 4y = 36 10y = 5 system of equations problem

OpenStudy (anonymous):

2x + 4y = 36 10y = 5 10y = 5 y = 5/10 = 1/2 Use the value of y in equation one: 2x + 4(1/2) = 36 2x + 2 = 36 2x = 36 - 2 2x = 34 x = 34/2 = 17